OpenStudy (anonymous):

what is the process or step you need to follow when adding and subtracting radical expressions?

OpenStudy (anonymous):

really depends on the radical expression. an example would be good

OpenStudy (anonymous):

If they are like terms, collect them and try to factor them.

OpenStudy (anonymous):

Can some one explain me the process with an example.

OpenStudy (anonymous):

\[\sqrt{32}+\sqrt{8}\] \[\sqrt{16*2}+\sqrt{4*2}\] \[\sqrt{16}*\sqrt{2}+\sqrt{4}*\sqrt{2}\] \[4\sqrt{2}+2\sqrt{2}\] \[6\sqrt{2} = \sqrt{72}\]

OpenStudy (anonymous):

actually since you asked the question it would be best if you provided the example

OpenStudy (anonymous):

Well, I need that someone explain me the process. That's why I ask.

OpenStudy (anonymous):

Did you not understand the process given above ?

OpenStudy (anonymous):

yes I did but I understand better when I can see the process in a problem

OpenStudy (anonymous):

Which part did you not understand ?

OpenStudy (anonymous):

Abdul you help me yesterday and I had to log off without answer you what I didn't understand in the problem. Can you explain me now?

OpenStudy (anonymous):

Which part did you not understand ?